# Convexity on manifolds without focal points and spectrum of the Laplacian

### v1

@@ 4. Spectral theoretic applications
for $\lambda\in{\mathbb{C}}$ and $v\in S_{o}M$. Here $\phi_{\lambda}$ is the unique function on $[0,\infty)$ with $\phi_{\lambda}(0)=1$ and satisfying $\Delta(\phi_{\lambda}\circ d_{o})=(\lambda^{2}+\frac{h^{2}}{4})(\phi_{\lambda}\circ d_{o})$, where $d_{o}$ denotes the distance from $o$. Let ${\rm{L}}_{o}^{2}(M)$ denote the closed subspace of ${\rm{L}}^{2}(M)$ consisting of those functions which are radial around $o$ in ${\rm{L}}^{2}(M)$. It is known from [3, Theorem 4.7] that the ${\mathcal{F}}^{o}$ is an unitary operator from ${\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$. Here the $c$-function coincides with the Harish Chandra $c$-function on $Im(\lambda)\leq 0$, for rank one symmetric spaces of noncompact type.
We recall that for a real-valued measurable function $f$ on a $\sigma-$finite measure space $(X,\mu)$, the multiplication operator $M_{f}$ on ${\rm{L}}^{2}(X,\mu)$ is self-adjoint with domain $\mathcal{D}(M_{f}):=\{g\in{\rm{L}}^{2}(X,\mu):M_{f}g\in{\rm{L}}^{2}(X,\mu)\}$. Let $m$ be the function on $[0,\infty)$ defined by $m(\lambda)=\lambda^{2}+\frac{h^{2}}{4}$. Then we see that the multiplication operator $M_{m}$ on ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$ is self-adjoint on
@@
Moreover, the spectrum of a multiplication operator $M_{m}$ is equal to the essential range of $m$ [4, Theorem 4.5]. Here
$ess-range(m):=\{x\in{\mathbb{R}}:\mu(m^{-1}(B(x,\epsilon)))>0~{}\forall~{}\epsilon>0\}.$
Now to determine the spectrum of $M_{m}$ we argue as follows. For any $x\in{\mathbb{R}}$ and $\epsilon>0$,
$m^{-1}(B(x,\epsilon))=\{\lambda\in[0,\infty):x-\epsilon<\lambda^{2}+\frac{h^{2}}{4}<x+\epsilon\}$
Note that when $x<\frac{h^{2}}{4}$ and $\epsilon=-(x-\frac{h^{2}}{4})$, $m^{-1}(B(x,\epsilon))$ is empty and so $x\notin ess-range(m)$. Equivalently, $ess-range(m)\subseteq[\frac{h^{2}}{4},\infty)$. Now for any $x\in[\frac{h^{2}}{4},\infty)$ and $\epsilon>0$ we have,
$m^{-1}(B(x,\epsilon))=\begin{cases}\left[0,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)&\text{if}~{}x-\frac{h^{2}}{4}-\epsilon<0\\ \\ \left[\sqrt{x-\frac{h^{2}}{4}-\epsilon},\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)&\text{if}~{}x-\frac{h^{2}}{4}-\epsilon\geq 0.\\ \end{cases}$
Using the estimates on $|c(\lambda)|^{-1}$ given in [3, §4.1] we see that, for $C>0$ and $K=\sqrt{x-\frac{h^{2}}{4}+\epsilon}>0$,
$\int\limits_{m^{-1}(B(x,\epsilon))}C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda\geq\frac{C_{0}}{C}\int\limits_{m^{-1}(B(x,\epsilon))}|\lambda|^{2}{\rm{d}}\lambda>0.$
It follows that $ess-range(m)=[\frac{h^{2}}{4},\infty)$ and so the spectrum of $M_{m}$ is equal to $[\frac{h^{2}}{4},\infty)$. For functions that are not necessarily radial, a notion of Fourier transform, relative to a point, was introduced in [3, §5] where it was shown to be an isometry but the unitarity is not known. This Fourier transform coincides with the spherical Fourier transform ${\mathcal{F}}^{o}$ for radial functions [3, Proposition 5.3]. Then by using [5, Lemma 6.5], we see that
@@
The spectrum of the Laplacian $\Delta$ on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ is equal to $[h^{2}/4,\infty)$.
@@
$[h^{2}/4,\infty)\subseteq\sigma(\Delta)$ where $\sigma(\Delta)$ is the spectrum of $\Delta$ with domain $\mathcal{D}(\Delta)$.
@@
and by uniqueness $R_{o}g=g$ i.e., $\Delta-\mu$ is surjective from $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}_{o}^{2}(M)$ and hence invertible. Since $(\Delta-\mu)^{-1}$ is bounded on ${\rm{L}}^{2}(M)$ and ${\rm{L}}_{o}^{2}(M)$ is a closed subspace, $(\Delta-\mu)^{-1}$ is also bounded on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$. The claim now follows from Lemma 4.5. ∎
The equality of the spectrum now follows from Corollary 4.4 and Lemma 4.6. Therefore, we have the following theorem.
Theorem 4.7.
Let $M$ be a simply connected noncompact nonflat harmonic manifold without focal points. Then the spectrum of Laplacian is purely absolutely continuous and is equal to $[h^{2}/4,\infty)$.
[…]
@@
From the structure theory for noncompact connected real semisimple Lie algebras, we have the set of all positive roots denoted by $\sum^{+}$. Let $m_{\alpha}$ denote the multiplicity of $\alpha\in\sum^{+}$. The finite reflection group $W$ generated by the roots is known as the Weyl group. We denote the tangent vector that is dual to $2\rho=\sum\limits_{\alpha\in\sum^{+}}m_{\alpha}\alpha$ also by $2\rho$. The arguments in [27, Lemma 1] show that for each $v\in S_{o}M$, $\Delta b_{v}$ is the constant function given by $\sum\limits_{i}\sqrt{\lambda_{i}(v)}={\langle v},{2\rho\rangle}$. This constant is the mean curvature of horospheres with respect to $b_{v}$. Here we argue as in Proposition 4.1 by using Stoke’s theorem to get a lower bound for Cheeger’s constant namely, ${\langle v},{2\rho\rangle}\leq h_{Cheeger}$ for all $v\in S_{o}M$. In particular, for $v=\frac{\rho}{\left\lVert\rho\right\rVert}$ we have $\left\lVert\rho\right\rVert\leq h_{Cheeger}$ and hence by Cheeger’s inequality we obtain
(4.7) $\left\lVert\rho\right\rVert^{2}\leq\lambda_{0}(M).$
Observe that $M$ has negative Ricci curvature [12, §2.14, p. 96] and has no focal points. With notations as in Theorem 3.1 we have,
@@
The absolute continuity of the spectrum $\sigma(\Delta)$ is then established by invoking [28, Theorem 2]. Further, using (4.7) we have,
Theorem 4.8.
Let $M$ be a symmetric space of noncompact type. Then the spectrum of the Laplacian on $M$ is absolutely continuous and is equal to $[\left\lVert\rho^{2}\right\rVert,\infty)$.
Here we use the spherical Fourier transform for the $K$-bi-invariant functions on $G$ to prove the equality of the spectrum. The radial part of the Laplacian $\Delta$ for the action of $K$ on $G/K$ is given by [19, §4.5, p. 427],
(4.8) $\Delta_{rad}\phi_{\lambda}=({\langle\lambda},{\lambda\rangle}+{\langle\rho},{\rho\rangle})\phi_{\lambda},\quad\lambda\in\mathfrak{a}_{{\mathbb{C}}}^{\ast}$
where $\mathfrak{a}_{{\mathbb{C}}}^{\ast}$ the complexification of the dual $\mathfrak{a}^{\ast}$ of $\mathfrak{a}$ and $\phi_{\lambda}(gK)=\displaystyle{\int\limits_{K}e^{(i\lambda+\rho)A(kg)}~{}{\rm{d}}k}$ is the unique spherical function on $G$ with the normalization $\phi_{\lambda}(eK)=1$ and $A$ is referred to as the Iwasawa projection $A:G\to{\mathfrak{a}}$. Two spherical functions are identical up to the action by the Weyl group $W$ on $\mathfrak{a}_{{\mathbb{C}}}^{\ast}$. Let us denote the space of smooth compactly supported $K-$invariant functions on $G/K$ by $C_{c}^{\infty}(K\backslash G/K)$ and its closure in ${\rm{L}}^{2}(G/K)$ by ${\rm{L}}_{rad}^{2}(G/K)$. Using the inversion formula for the spherical transform [19, Chapter 4, Theorem 7.5] the map $f\mapsto\widehat{f}$
$\widehat{f}(\lambda)=\int\limits_{G/K}f(gK)\phi_{\lambda}(gK),\quad f\in C_{c}^{\infty}(K\backslash G/K)$
[…]
(4.9) $\int\limits_{m^{-1}(B(x,\epsilon))}|\textbf{c}(\lambda)|^{-2}{\rm{d}}\lambda>0.$
for $x\geq\left\lVert\rho\right\rVert^{2}$ and $\epsilon>0$. Note that $|\textbf{c}(\lambda)|^{2}$ defined on ${\mathfrak{a^{\ast}}}$ is invariant under the action of $W$ i.e., $|\textbf{c}(w\lambda)|^{2}=|\textbf{c}(\lambda)|^{2}$ for all $w\in W$ [19, §IV.7, Equation 16] and hence descends to a function on ${\mathfrak{a}^{\ast}}/W$. Since $\lambda\mapsto\left\lVert\lambda\right\rVert^{2}$ is continuous and
$m^{-1}(B(x,\epsilon))=\left\{\lambda\in{\mathfrak{a}^{\ast}}/W:x-\epsilon-\left\lVert\rho\right\rVert^{2}<\left\lVert\lambda\right\rVert^{2}<x+\epsilon-\left\lVert\rho^{2}\right\rVert\right\}$
we see that $m^{-1}(B(x,\epsilon))$ contains an open ball in $\mathfrak{a}^{\ast}/W$. Using [15, Chapter 3, p. 101, Theorem 6.3.4], we have that the map $\lambda\mapsto|\textbf{c}(\lambda)|^{-2}$ is analytic on $\mathfrak{a}^{\ast}$ and thereby analytic on $\mathfrak{a}^{\ast}/W$. Thus, (4.9) holds for $x\geq\left\lVert\rho\right\rVert^{2}$ and $\epsilon>0$. This, in particular, shows that the spectrum of $\Delta_{rad}$ is equal to $[\left\lVert\rho\right\rVert^{2},\infty)$. We argue as in Lemma 4.6 to show the equality in Theorem 4.8.

### v2

@@ 4. A construction using Busemann functions and its applications
for $\lambda\in{\mathbb{C}}$ and $v\in S_{o}M$. Here $\phi_{\lambda}$ is the unique function on $0,\infty)$ with $\phi_{\lambda}(0)=1$ and satisfying $\Delta(\phi_{\lambda}\circ d_{o})=(\lambda^{2}+\frac{h^{2}}{4})(\phi_{\lambda}\circ d_{o})$, where $d_{o}$ denotes the distance from $o$. Let ${\rm{L}}_{o}^{2}(M)$ denote the closed subspace of ${\rm{L}}^{2}(M)$ consisting of those functions which are radial about $o$ in ${\rm{L}}^{2}(M)$. It is known from [BKP21, Theorem 4.7] that ${\mathcal{F}}^{o}$ is a unitary operator from ${\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$. Here the $c$-function coincides with the Harish Chandra $c$-function on $Im(\lambda)\leq 0$, for rank one symmetric spaces of noncompact type.
We recall that for a real-valued measurable function $f$ on a $\sigma$-finite measure space $(X,\mu)$, the multiplication operator $M_{f}$ on ${\rm{L}}^{2}(X,\mu)$ is self-adjoint with domain $\mathcal{D}(M_{f}):=\{g\in{\rm{L}}^{2}(X,\mu):M_{f}g\in{\rm{L}}^{2}(X,\mu)\}$. Let $m$ be the function on $0,\infty)$ defined by $m(\lambda)=\lambda^{2}+\frac{h^{2}}{4}$. Then we see that the multiplication operator $M_{m}$ on ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$ is self-adjoint on
@@
Moreover, the spectrum of a multiplication operator $M_{m}$ is equal to the essential range of $m$ [Bor20, Theorem 4.5]. Here
$ess-range(m):=\{x\in{\mathbb{R}}:|m^{-1}(x-\epsilon,x+\epsilon))|>0,\penalty 10000\ \forall\penalty 10000\ \epsilon>0\}.$
where $|\cdot|$ denotes the measure. Now to determine the spectrum of $M_{m}$ we argue as follows. For any $x\in{\mathbb{R}}$ and $\epsilon>0$,
$m^{-1}(x-\epsilon,x+\epsilon)=\{\lambda\in[0,\infty):x-\epsilon<\lambda^{2}+\frac{h^{2}}{4}<x+\epsilon\}$
Note that when $x<\frac{h^{2}}{4}$ and $\epsilon=-(x-\frac{h^{2}}{4})$, $m^{-1}(x-\epsilon,x+\epsilon)$ is empty and so $x\notin ess-range(m)$. Equivalently, $ess-range(m)\subseteq[\frac{h^{2}}{4},\infty)$. Now for any $x\in[\frac{h^{2}}{4},\infty)$ and $\epsilon>0$ we have,
[…]
We recall the estimates for $|c(\lambda)|^{-1}$ given in [BKP21, §4.1].
If $x-\frac{h^{2}}{4}-\epsilon<0$, then $\lambda\in\left[0,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)$. If $\sqrt{x-\frac{h^{2}}{4}+\epsilon}\leq K$ then we have
$\frac{1}{C}\lambda\leq|c(\lambda)|^{-1}$
and so
$\int\limits_{m^{-1}(x-\epsilon,x+\epsilon)}C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda\geq\frac{C_{0}}{C}\int\limits_{0}^{\sqrt{x-\frac{h^{2}}{4}+\epsilon}}\lambda^{2}{\rm{d}}\lambda>0,\quad\text{since}\penalty 10000\ x-\frac{h^{2}}{4}+\epsilon>0$
Now, if $K\in\left(0,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)$ then
$\begin{cases}\frac{1}{C}\lambda\leq|c(\lambda)|^{-1},&\lambda\in[0,K)\\ \frac{1}{C}\lambda^{\frac{n-1}{2}}\leq|c(\lambda)|^{-1},&\lambda\in\left[K,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)\end{cases}$
holds thereby
$\int\limits_{m^{-1}(x-\epsilon,x+\epsilon)}C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda\geq\begin{cases}\frac{C_{0}}{C}\int\limits_{0}^{K}\lambda^{2}{\rm{d}}\lambda,&\lambda\in[0,K)\\ \\ \frac{C_{0}}{C}\int\limits_{K}^{\sqrt{x-\frac{h^{2}}{4}+\epsilon}}\lambda^{n-1}{\rm{d}}\lambda,&\lambda\in\left[K,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)\end{cases}$
and hence the integral is always strictly positive. A similar argument shows that the integral remains strictly positive for the other case in (13). It follows that $ess-range(m)=[\frac{h^{2}}{4},\infty)$ and so the spectrum of $M_{m}$ is equal to $\frac{h^{2}}{4},\infty)$. For functions that are not necessarily radial, a notion of Fourier transform, relative to a point, was introduced in [BKP21, §5] where it was shown to be an isometry but the unitarity is not known. This Fourier transform coincides with the spherical Fourier transform ${\mathcal{F}}^{o}$ for radial functions [BKP21, Proposition 5.3]. Then we see as in [Bra24, Lemma 6.5], that
@@
The spectrum of the Laplacian $\Delta$ on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ is equal to $h^{2}/4,\infty)$.
@@
$h^{2}/4,\infty)\subseteq\sigma(\Delta)$ where $\sigma(\Delta)$ is the spectrum of $\Delta$ with domain $\mathcal{D}(\Delta)$.
@@
and since $\Delta-\mu$ is invertible we have $R_{o}g=g$ i.e., $\Delta-\mu$ is surjective from $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}_{o}^{2}(M)$. Since $(\Delta-\mu)^{-1}$ is bounded on ${\rm{L}}^{2}(M)$ and ${\rm{L}}_{o}^{2}(M)$ is a closed subspace, $(\Delta-\mu)^{-1}$ is also bounded on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$. The claim now follows from Lemma 4.9. ∎
The equality of the spectrum is obtained from Corollary 4.7 and Lemma 4.10. In summary, we have the following proposition.
Proposition 4.11.
Let $M$ be a simply connected noncompact (nonflat) harmonic manifold without focal points. Then the spectrum of Laplacian is purely absolutely continuous and is equal to $h^{2}/4,\infty)$.
As remarked before, rank-one symmetric spaces of noncompact type, Damek-Ricci spaces are among the classes of manifolds which satisfy the hypothesis of Proposition 4.11. In fact, it is known that an irreducible symmetric space is harmonic if and only if it is rank one [Kni16]. We shall see that, on the other hand, Riemannian symmetric spaces of noncompact type and higher rank are not even asymptotically harmonic and therefore Theorem 4.3 does not apply. Nonetheless, we will indicate below that our method can still be employed to show an analogue of Proposition 4.11 in this setting. Let $M$ be an $n$-dimensional symmetric space of noncompact type i.e., $M$ is a simply connected Riemannian manifold with parallel curvature tensor $R$ [Ebe96, Chapter 2]. Note that $M$ is necessarily of the form $G/K$ where $G=NAK$ is the Iwasawa decomposition of a noncompact connected semisimple Lie group $G$ with finite center. The rank of the symmetric space is the dimension $d$ of $\mathfrak{a}$, the Lie algebra of $A$. The Killing form on the Lie algebra of $G$ induces an inner product on $\mathfrak{a}$ denoted by $\langle\cdot$$\cdot\rangle$}. Let us fix a base point $o=eK\in G/K$. If the rank of $M$ is atleast two, then its sectional curvature vanishes on the $G$-conjugates of $A.o$ in $G/K$. These are the maximal totally geodesic flat submanifolds in $G/K$. For $v\in S_{o}M$, consider the geodesic $\gamma_{v}$ satisfying $\gamma_{v}^{{}^{\prime}}(0)=v$ and let $R_{v}$ be the curvature tensor restricted to $\gamma_{v}$. Choose an orthonormal basis $\{e_{1},\cdots,e_{n-1}\}$ of eigenvectors of $R_{v}$ for the subspace in $T_{o}M$ orthogonal to $v$ i.e.,
@@
From the structure theory for noncompact connected real semisimple Lie algebras, we have the set of all positive roots denoted by $\sum^{+}$. Let $m_{\alpha}$ denote the multiplicity of $\alpha\in\sum^{+}$. The finite reflection group $W$ generated by the roots is known as the Weyl group. We denote the tangent vector that is dual to $2\rho=\sum\limits_{\alpha\in\sum^{+}}m_{\alpha}\alpha$ also by $2\rho$. The arguments in [Wan15, Lemma 1] show that for each $v\in S_{o}M$, $\Delta b_{v}$ is the constant function given by $\sum\limits_{i}\sqrt{\lambda_{i}(v)}={\langle v},{2\rho\rangle}$. This constant is the mean curvature of horospheres with respect to $b_{v}$. Here we argue as in Proposition 4.4 by using Stoke’s theorem to get a lower bound for Cheeger’s constant namely, ${\langle v},{2\rho\rangle}\leq h_{Cheeger}$ for all $v\in S_{o}M$. In particular, for $v=\frac{\rho}{\left\lVert\rho\right\rVert}$ we have $\left\lVert\rho\right\rVert\leq h_{Cheeger}$ and hence by Cheeger’s inequality we obtain
(14) $\left\lVert\rho\right\rVert^{2}\leq\lambda_{0}(M).$
Observe from [Wan15], that along a geodesic $\gamma_{v}$ the Jacobi field $Y_{i}$ satisfying the initial condition $Y_{i}(0)=0$ and $Y_{i}^{{}^{\prime}}(0)=e_{i}$ is given by the explicit formula
$Y_{i}(t)=\frac{1}{\sqrt{\lambda_{i}(v)}}\sinh\left(\sqrt{\lambda_{i}(v)}t\right)E_{i}(t)$
where $E_{i}$ is a parallel orthonormal frame field along $\gamma_{v}$ with $E_{i}(0)=e_{i}$. Then we see that
$Y(t)=\sum\limits_{i=1}^{n}{\langle w},{e_{i}\rangle}Y_{i}(t)$
solves the Jacobi equation (1) with initial conditions $Y(0)=0$ and $Y^{{}^{\prime}}(0)=w$ for arbitrary $w\in SM$. That $\left\lVert Y(t)\right\rVert$ is strictly increasing can then be deduced from the fact that $\sinh$ is an increasing function and so, it follows from [O’S74, Proposition 4] that $M$ has no focal points. It is known that $M$ is also Einstein with negative Ricci curvature [Ebe96, §2.14, p. 96]. Therefore Theorem 4.1 applies and we have,
@@
As before, the absolute continuity of the spectrum $\sigma(\Delta)$ is established by using [Xav88, Theorem 2]. That the spectrum is equal to $\left\lVert\rho^{2}\right\rVert,\infty)$ can be shown analogous to the case of harmonic manifolds by using the spherical Fourier transform and estimates on the c-function found in [GV88, Chapter 3, p. 101, Theorem 6.3.4].